# Optimizing Allreduce Operations for Heterogeneous Architectures with Multiple Processes per GPU

### v1

@@ 4.1. MPI Analysis
Figure 10. The multi-lane all-reduce with multiple processes per GPU.
((a)) Tuolumne (RCCL)
((b)) DeltaAI (NCCL)
Figure 11. Standard CCL algorithm with a size of $2^{29}$ floats, varying node counts
@@
Finally, Figure 8 compares the performance of all algorithms, including where the ring algorithm is used in the inter-node stage of the multi-lane all-reduce, comparing $1$ versus $16$ processes per GPU. The multi-lane algorithm with $16$ processes per GPU achieves up to 2.42.4x speedup over the standard MPI_Allreduce for large data sizes.
4.2. CCL Analysis
This section analyzes the impact of using multiple processes per GPU across CCL all-reduce operations. The performance of NCCL is analyzed on a state-of-the-art NVIDIA Grace Hopper architecture, Delta AI, using CUDA 12.3 and NCCL 2.19.3. The performance of RCCL is analyzed on one of the fastest AMD architectures, Tuolumne, using ROCm 6.4.0 and RCCL 2.22.3. Each benchmark was run 3 separate times; error bars in plots show run to run variation.
Each process performs an independent CCL all-reduce operation on their assigned subset of the buffer. As CCL all-reduce operations take a stream, each process per GPU performs this reduction on a separate stream to allow increased parallelism. All processes were assigned to the GPU adjacent to their NUMA region.
@@ 4.2. CCL Analysis
Figure 11 shows the performance of a CCL all-reduce at the largest size when varying the number of processes per GPU from 1 to 16 on Tuolumne and DeltaAI. On the left featuring Tuolume using RCCL, we see steady speedup as the number of nodes increases with 2 and 4 processes per GPU, yielding speedups of 1.711.71x and 1.591.59x at 8 nodes, respectively. These speedups begin to decrease at large scales. Using more than 4 processes per GPU does not show any benefit on Tuolumne past 8 nodes with 8 and 16 processes per GPU showing significant slowdown.
On the right featuring DeltaAI using NCCL, we see substantial speedup at all node counts with 16 processes per GPU, yielding the most significant speedup of 1.771.77x at 16 nodes. 4 and 8 processes per GPU show speedup only at small scales, while using 2 processes per GPU starts to show benefit at 16 nodes.
There are many potential reasons for this speedup, including improved network utilization and improved use of NCCL channels. To determine if NCCL channels were impacted, the environment variables NCCL_MIN_NCHANNELS and NCCL_MAX_NCHANNELS were set to 4 in order to eliminate the case of low GPU core utilization during memory operations; no differences were observed in timings.
5. Conclusions and Future Work
@@
In this paper, we proposed to utilize more of the available CPU cores across all GPUs on a node to speed up the GPU-aware all-reduce collective. Using an Ampere-based system (Delta) on OpenMPI with our standard and lane-aware approaches, we showed up to an 2.452.45x speedup on large buffers. Using RCCL with our standard approach on an El Capitan-based system, we showed a 1.711.71x speedup. Using NCCL on a Hopper-based system, we showed a 1.771.77x speedup. We conclude that using more of the available CPU cores on modern systems should enhance all-reduce performance at large scales.
@@
Emerging systems like Tuolumne feature memory architectures that are shared between the CPU and GPU. Future work could potentially examine tradeoffs using CPUs for communication and GPUs for computation, allowing these steps to overlap while minimizing synchronization costs.
@@ Acknowledgments
This research used the Tuolumne supercomputer at Lawrence Livermore National Laboratory. This research used both the DeltaAI advanced computing and data resource, which is supported by the National Science Foundation (award OAC 2320345) and the State of Illinois, and the Delta advanced computing and data resource which is supported by the National Science Foundation (award OAC 2005572) and the State of Illinois. Delta and DeltaAI are joint efforts of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications.
Generative AI was used to improve the formatting of data within plots.

### v2

@@ V-A Host Copy Analysis
Figure 12: Standard and multi-lane algorithms with multiple processes per GPU on 8 nodes of Delta, varying buffer sizes
Figure 13: Standard and multi-lane algorithms with multiple processes per GPU of $2^{27}$ floats on Delta, varying node counts
@@
Finally, Figures 12 and 13 compare the performance of all algorithms, including where the ring algorithm is used in the inter-node stage of the multi-lane all-reduce, comparing $1$ versus $16$ processes per GPU. The multi-lane algorithm with $16$ processes per GPU achieves up to 2.42.4x speedup over the standard MPI_Allreduce for large data sizes. These results agree with our ping-pong microbenchmark on Delta.
V-B GPUDirect RDMA Analysis
The performance of GPUDirect RDMA-enabled MPI is analyzed on one of the fastest AMD architectures, Tuolumne, using Cray MPICH 9.0.1 and ROCm 6.4.0. For benchmarks using more than one process per GPU, data to be communicated was copied to and from an allocated device buffer before and after MPI communication. This is because Cray MPICH does not support buffers shared via ROCm IPC. Each process performs an independent all-reduce operation on their assigned subset of the buffer. All processes were assigned to the GPU adjacent to their NUMA region. Each benchmark was run 3 separate times; error bars in plots show run to run variation.
Figure 14: Standard algorithm with multiple processes per GPU on 32 nodes of Tuolumne, varying buffer sizes
Figure 15: Standard algorithm with multiple processes per GPU of $2^{29}$ floats on Tuolumne, varying node counts
Figures 14 and 15 show the performance of an all-reduce when varying the number of processes per GPU from 1 to 4 on Tuolumne with each MI300A in SPX mode. Figure 14 shows timings on 32 nodes at different sizes of the communication buffer where using multiple processes per GPU shows speedup for the largest buffer sizes.
@@ V-B GPUDirect RDMA Analysis
Figure 15 shows the timings of the largest buffer size as the number of nodes increases. We see steady speedup as the number of nodes increases with 2 and 4 processes per GPU, yielding speedups of 1.171.17x and 1.11.1x at 32 nodes, respectively. These speedups begin to decrease at large scales; more significant speedups of up to 1.211.21x are seen at 8 nodes. Using more than 4 processes per GPU does not show any benefit on Tuolumne.
It is noted that a small benchmark was performed on 8 nodes of Tuolumne with the System Direct Memory Access (SDMA) engines of the MI300A both enabled and disabled. Consistent with previous findings regarding small-scale inter-APU communication [36], the use of SDMA engines does not appear to impact the advantages provided by using multiple processes per GPU.
V-B1 MI300A Partitioning Modes
Figure 16: Standard and multi-lane algorithms with multiple processes per GPU on 32 nodes across the SPX/TPX/CPX modes of Tuolumne, varying buffer sizes
Figure 17: Standard and multi-lane algorithms with multiple processes per GPU of $2^{28}$ floats across the SPX/TPX/CPX modes of Tuolumne, varying node counts
The AMD MI300A APUs are also tested in the TPX and CPX partitioning modes. Figures 16 and 17 show the speedups of using two processes per GPU for both proposed algorithms against the standard MPI_Allreduce implementation for each partitioning mode  [^1] It is important to note that speedups from each partitioning mode cannot be directly compared since the buffer size per logical GPU is not scaled with the number of logical GPUs.. Figure 16 shows speedups on 32 nodes of Tuolumne with different buffer size. In SPX and CPX modes with either the standand or multi-lane algorithm, respectively, there is a significant advantage to using two processes per GPU for larger buffers.
Figure 17 shows speedups of both algorithms while the number of nodes increases at the largest buffer size tested across all partitioning modes. One of the novel algorithms presented provides speedup in SPX and CPX modes; we see that the multi-lane multi-process algorithm yields speedups of up to 33x over the standard algorithm in CPX mode while the standard multi-process algorithm shows no speedup. Performance in TPX mode does not show benefit at scale for either algorithm. The standard multi-process algorithm is only advantageous in SPX mode. Thus, the multi-lane algorithm benefits from more logical GPUs potentially due to the increased memory locality near each XCD die [3] with smaller communication sizes from the initial call to reduce-scatter. The standard algorithm exhibits the opposite, potentially benefiting from less logical GPUs per node, in agreement with the SPX ping-pong microbenchmark on Tuolumne.
We also observe that our multi-lane algorithm using two processes per GPU in CPX mode scales significantly better than the standard algorithm in any of the partitioning modes; the speedup of our multi-lane algorithm continues to increase as the number of nodes increases.
Figure 18: The multi-lane all-reduce with multiple processes per GPU on the MI300A partitioning modes on Tuolumne
The performance of the multi-lane algorithm as the number of nodes increases can be seen in Figure 18. We observe that using multiple processes per GPU benefits the multi-lane algorithm in SPX and CPX modes, yielding 1.51.5x and 2.12.1x speedup, respectively, emphasizing that the speedup of the multi-lane algorithm in Figure 17 is due to both the algorithmic structure, the use of multiple processes per GPU, and the fine-grained parallelism offered by CPX partitioning.
VI Conclusions and Future Work
@@
In this paper, we proposed to utilize more of the available CPU cores across all GPUs on a node to speed up the GPU-aware all-reduce collective. Using an Ampere-based system (Delta) on OpenMPI backed by host copy communication with our standard and lane-aware approaches, we showed up to an 2.452.45x speedup on large buffers. Using Cray MPICH backed by GPUDirect RDMA communication with our standard approach on an El Capitan-based system (Tuolumne), we showed a 1.171.17x speedup. Across the three partitioning modes of the AMD MI300A on Tuolumne, we showed that the multi-process multi-lane all-reduce is highly performant in CPX mode, yielding a 33x speedup at scale.
@@
Emerging systems like Tuolumne feature memory architectures that are shared between the CPU and GPU. Future work could potentially examine tradeoffs in scientific applications by driving communication with more CPU cores while using GPUs for local computation, allowing these steps to overlap while minimizing synchronization costs.
@@ Acknowledgments
This research used the Tuolumne supercomputer at Lawrence Livermore National Laboratory. This research used the Delta advanced computing and data resource which is supported by the National Science Foundation (award OAC 2005572) and the State of Illinois. Delta is a joint effort of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications.
Generative AI was used to improve formatting of plots.